40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3117     Entry time: Thu Jun 24 18:47:26 2010
Author: Frank 
Type: DAQ 
Category: IOO 
Subject: VME crate rebooted  

we had to reboot the IOO VME crate right before lunch as the DAQ wasn't working correct meaning showing no real signals anymore, only strange noise. The framebuilder and everything else was working fine at that time.

  • The channel used for the phase noise measurement stopped showing any useful signal right after midnight, so all the other IOO-MC signals.
  • The data taken with those channels showed something like a 140 counts or so of steady offset with something which looked like the last bit fluctuating.
  • Whatever signal we connected to the input it didn't change at all, floating/shorted input, sine wave etc.
  • the other channels for the MC which we checked showed the same strange behaviour

As the other channels showed the same effect we decided to reboot the crate and everything was fine afterwards.

ELOG V3.1.3-